Shift report 20000802 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Aug 02 2000 - 16:43:56 EDT

  • Next message: http account: "Shift report 20000803 16:00-24:00"

    ------------------------------------------------------------
    Date: Wed Aug  2 16:43:55 2000, Shift: 08:00-16:00
    Supervisor: DR, with: TF & VS
    Log book#13 pages 129 - 129
    
    Shift Summary:
    ------------------------------------------------------------
    1. "working on injection" 
    2. none
    3. FFS1440 not responding, BBL4032 goes into "unknown" state
       after 10-30 min. due to tripping channels. Same 
       difficulties getting the magnets on as during 
       last evening shift.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Aug 02 2000 - 16:50:53 EDT